A spring has a spring constant of 1000N/m. It is compressed by 15cm, then launches a 200g block on a frictionless horizontal surface, it then reaches a 30 DEGREE incline and the blocks coefficient of kinetic friction on incline is 0,20.

What distance 'd' does the block sail through the air?

Please assist.

{ ___d__}
/
[ ] ____/

some type of representation.
1
Expert's answer
2020-05-07T10:00:09-0400

As per the given question,

Spring constant"(k)= 1000N\/m."

Compression distance "(x)=15cm"

Mass of the object "(m)=200g=0.2 kg"

Angle of inclination of the plane "(\theta)=30^\circ"

Coefficient of kinetic friction "(\mu)=0.2"

let l is the distance traveled on the inclined plane

Applying the conservation of energy,

"\Rightarrow \dfrac{kx^2}{2}=mg\sin\theta \times l"

"\Rightarrow l =\dfrac{kx^2}{2mg\sin\theta}=\dfrac{1000\times 0.15\times 0.15}{2\times 0.2 \times 9.8 \times \sin 30}"

"\Rightarrow l =\dfrac{225}{1.96}=114.8m"

and height raised by block =57.4m"=l\sin\theta =57.4m"